OpenStudy (anonymous):

The radius of a 10 inch right circular cylinder is measured to be 4 inches, but with a possible error of ±0.1 inch. What is the resulting possible error in the volume of the cylinder? Include units in your answer.

OpenStudy (anonymous):

@freckles @Luigi0210 @Nnesha @Compassionate can u plz help? :)

OpenStudy (anonymous):

@iambatman

OpenStudy (freckles):

\[\Delta V \approx \Delta x \cdot \frac{dV}{dx}|_{x=a} \] So we are given \[\Delta x =\pm 0.1 \\ V=\pi r^2 h \] \[\text{ We want to find } \Delta V \] Hmmm I wonder if both r and h are changing...

OpenStudy (freckles):

It sounds like the radius is the one with the possible error right?

OpenStudy (freckles):

So we will assume h is constant and take derivative of V with respect to r And the equation I have above we will replace x with r instead

OpenStudy (freckles):

\[V=\pi h r^2 \] So I assume you can find the d (r^2)/dr

OpenStudy (freckles):

\[\frac{dV}{dr}=\frac{d( \pi h r^2)}{dr}= \pi h \frac{d(r^2)}{dr} \\ \frac{dV}{dr}=\pi h \frac{d (r^2)}{dr}\]